# CONSENT: A Negotiation Framework for Leveraging User Flexibility in Vehicle-to-Building Charging under Uncertainty

### v1

@@ Abstract.
1. Introduction
@@
The rapid growth of electric vehicles (EVs), fueled by sustainability goals and advances in battery technology (al2019review; wu2023electric), presents a significant challenge for energy management. When EV charging is uncoordinated, it can trigger sharp demand spikes, inflate energy costs, and threaten grid stability (liu2021ev). This has spurred research into coordinated charging, particularly in the vehicle-to-building (V2B) context where operators can leverage parked EVs for peak shaving (wang2018ev). However, creating an effective V2B system requires navigating a set of deeply interconnected challenges that span both technical control and human behavior.
@@
At the core of the control problem lies the interplay between uncertainty and long-term rewards (liu2025). A charging controller must operate with incomplete knowledge of future building loads and EV arrivals, yet its decisions are judged against long-term objectives like minimizing the monthly peak demand charge, a sparse and delayed reward. This makes myopic, reactive control insufficient. This technical challenge is further complicated by the need for incentive alignment (bae2021inducing). The operator’s goal of deferring charging to minimize electricity costs, driven by both energy prices and monthly peak-power demand charges, directly conflicts with the user’s desire for convenience and a full battery (he2022optimal). A negotiation must therefore quantify the value of a user’s flexibility, but this value is itself a function of an uncertain future. Finally, any negotiation is vulnerable to strategic misreporting, where users may exaggerate their needs to extract larger incentives, undermining the system’s integrity. Addressing these issues requires more than just efficient scheduling; it demands a strategic bridge between the building’s stationary energy system and the dynamic mobility needs of its users. While prior work addresses these challenges in isolation, a key frontier remains: the co-optimization of the real-time charging control policy with a strategic, user-facing negotiation policy. This requires a framework that both makes optimal control decisions and generates fair, personalized incentives. While methods like Reinforcement Learning (RL) have shown promise for handling uncertainty, the co-optimization task particularly benefits from approaches that can explicitly handle hard constraints and adapt to new scenarios without extensive re-training. For this reason, we posit that the explicit constraint handling and computational efficiency of Model Predictive Control (MPC) provide a more direct and scalable solution (shapiro2021lectures; saleem2024stochastic; solar_mpc). An MPC-based approach offers robust uncertainty handling through sampling, which we show is comparable to an RL baseline, establishing MPC as the ideal engine for our framework.
@@
Leveraging this MPC foundation, we introduce our negotiation framework, CONSENT, which directly addresses the shortcomings of prior work (liu2025; ghosh2017menu; bae2021inducing). To achieve incentive alignment, users are presented with tailored incentives for voluntary adjustments to their requested state-of-charge (SoC) or departure time, grounded in empirical survey data. To prevent strategic misreporting, our mechanism includes formal guarantees of strategy-proofness, budget feasibility, and voluntary participation. By integrating this mechanism and validating with real-world data, we demonstrate that operator savings can be aligned with high user participation and satisfaction. Our contributions include:
@@
Figure 1. Negotiation workflow: Upon arrival, users receive forecast-based offers. Accepted offers are then optimally scheduled to meet the negotiated SoC.
2. Related Research
@@
Research in EV charging systems has largely advanced along two parallel streams. The first focuses on the charging control problem, where the goal is to optimize charging schedules to minimize operator costs like peak demand (he2022optimal; christensen2025scoping). A wide array of methods have been developed, including rule-based heuristics and mathematical programming (nakahira2017smoothed; stankovic_EDF; tanguy2016optimization). More recent advances use Model Predictive Control (MPC) for real-time optimization, effectively adapting to dynamic loads, renewables, and prices (di2014electric; yang2023ev). However, this entire stream of research, including stochastic MPC (shapiro2021lectures), typically solves the problem from the operator’s perspective by assuming full system control, thereby overlooking the critical challenge of securing voluntary user participation (thran2025reserve; sen2025).
@@
Recognizing that user compliance cannot be guaranteed, the second stream of research focuses on the socio-technical challenge of user negotiation and cooperation. This work has produced incentive-based frameworks like menu-based pricing (ghosh2017menu) and iterative negotiation (wang2018ev; khan2022multi) to provide users with choices. These are often informed by behavioral models that predict user responses to inconvenience and rewards (shariatzadeh2025electric), with a key finding being the importance of user heterogeneity (andrenacci2023literature; zahringer2024watt). Despite being user-centric, these frameworks are often designed for large-scale aggregators and rarely integrated with the specific, real-time operational constraints of a single building.
@@
The limitations of these separate streams highlight a significant gap at their intersection: the co-optimization problem, which requires strategically coupling real-time control with empirically-grounded, user-centric negotiation. Tackling this frontier requires advanced methods capable of handling uncertainty and complex decisions. Reinforcement Learning (RL) has emerged as a powerful tool for learning control policies in such stochastic environments (liu2025; sen2025). However, for the specific co-optimization challenge, Model Predictive Control (MPC) presents a compelling alternative. While RL excels in exploration, the co-optimization problem benefits from MPC’s ability to explicitly handle hard constraints (e.g., user choices, charger limits) and efficiently evaluate hypothetical negotiation scenarios. By using Monte-Carlo sampling, MPC can robustly manage uncertainty, establishing it as a powerful engine for a framework that must bridge control and negotiation. Accordingly, our work addresses this co-optimization gap. We develop an MPC-based negotiation framework that operates directly on personalized user requests and is calibrated with empirical survey data, enabling voluntary and incentive-compatible cooperation at the building-user interface.
3. Problem Formulation
@@ 3.1. Specifications
(1) $\small\begin{split}&J^{\pi}(\bm{\theta}\mid\mathcal{W})=\sum_{t\in\mathcal{T}}g_{t}+K_{\mathbf{DC}}\cdot p^{\max}\\ &+K_{\mathbf{SOC}}\sum_{v\in\mathcal{V}}\left(e^{v}_{\mathrm{req}}-e^{v}_{\mathrm{dep}}\right)+K_{\mathbf{batt}}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q_{t}^{v}\end{split}$
@@
(2) $\small\begin{split}U^{\bm{\theta}_{\setminus v},\pi^{*}}(\theta_{v})=\mathbb{E}_{\bm{\hat{\theta}}_{\setminus v}\sim\psi}\Big[J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}\mid\mathcal{W}^{\prime})\\ -J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v}\mid\mathcal{W}^{\prime})\Big]\end{split}$
@@ 4. Approach
4.1. EV-Charger Assignment & Charging Policy
@@
4.2. Formulating the User Negotiation Options
@@
As EVs depart on schedule, chargers free up, and the system state updates to reflect the connected EVs and their SoCs, charger availability, building load, and any new arrivals, enabling accurate tracking of charging and user decisions. The SMDP state transition kernel $\mathcal{P}(s_{k+1},\Delta t_{k}\mid s_{k},a_{k};\pi^{ch})$ combines: (i) the stochastic EV inter-arrival distribution; and (ii) the deterministic evolution under charging policy $\pi^{ch}$ during $\Delta t_{k}$. We do not explicitly define the state-action transition function, as our simulator leverages data-driven generative models trained on actual operational records (10595624).
@@ 4.3. Negotiation Policy
(3) $\zeta^{v}=\sum_{t\in\mathcal{T}^{v}}g^{v}_{t}w_{t}-\alpha\cdot U({\theta}_{v})$
@@
Our negotiation engine selects incentive-backed charging offers for each arriving by integrating a Monte Carlo sampling based model predictive control (MC-MPC). Our approach belongs to the family of stochastic Model Predictive Control (SMPC) methods (MA2016). Algorithm 1 presents the user negotiation process.
@@
For each flexibility level $l$, the controller explores the corresponding continuous action subset $\mathcal{A}_{l}$ (from the SMDP) and solves a Mixed Integer Linear Program (MILP) implementing a Sample Average Approximation  (shapiro2021lectures). The MILP minimizes the average total cost over the sampled scenarios $\mathcal{F}$, yielding an optimal candidate action $a_{l}^{*}$ for each $l$:
(4) $a^{*}_{l}=\min_{\mathcal{A}_{l}}$ $\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\Bigg[\sum_{i=0}^{H(t)}g_{i,f}+K_{\mathrm{DC}}\cdot p^{\max}_{f}$
@@ 20 Update EV $v_{k}$’s requirements based on the negotiated choice $\theta_{k}$;
5. Survey and User Modeling
@@
User acceptance during negotiation is modeled via a logit-based choice function (bae2021inducing). The utility of acceptance for option $l$ by user type $n$ for charging cost $\zeta_{l}$ and given an external charging cost $\bar{E}$, is
(5) $\small Y^{n}_{l}=\Big[\bar{E}-(\zeta_{l}+I^{n}_{l})\Big]$
@@
Policy Description Building Cost (\$/month) User Cost (\$/kWh) Monthly User Cost (\$/month) Reject (%)
@@ 1 Building-Only Reference $8484\pm 2185$
2 Uncoordinated Charging (al2021slow) + Free Cost $9005\pm 2184$ $0.00$
3 Smart Charging + Free Cost (liu2025) $8785\pm 2144$ $0.00$
4 Smart charging + Fixed User Cost (svp) $8457\pm 2144$ $0.178$ $328\pm 61$
@@
6 Menu-based Negotiation (bae2021inducing; ghosh2017menu) $8479\pm 2153$ $0.157$ $309\pm 63$ 34.89
@@
Table 1. Comparative Evaluation of V2B Policies. The table progresses from a building-only reference to increasingly sophisticated baselines, culminating in our proposed negotiation framework (CONSENT). Lower values are better for all the metrics.
6. Experiments
We evaluate our negotiation framework using a modular simulator (10595624) of a building with multiple chargers, dynamic energy and demand charges, and realistic EV arrival and departure patterns. It captures stochastic user preferences, building load forecasts, and charger constraints, supporting online or ahead-of-time scheduling.
Hardware. Experiments ran on a 32-core, 6.2 GHz, 128 GB RAM machine using IBM CPLEX (ibm_cplex) for MILP.
Data Collection. We use data from a commercial site in Santa Clara, California (May 2023 – November 2024), including building power usage, charger activity, and vehicle telemetry. All EVs used support bidirectional charging, and charging is modeled with a three-segment piecewise linear SoC curve. We sampled $100$ such episodes, evenly split into $50$ for future sampling and $50$ for testing. Only weekdays are analyzed, as employee presence and demand charges are negligible on weekends. Figures 7 and 8 (Appendix F) show EV arrival/departure and SoC distributions. The site has controllable, 10 unidirectional (0–20 kW) and 5 bidirectional (±20 kW) chargers. Silicon Valley Power’s time-of-use and demand charge rates (svp) define $\mathcal{W}$ where peak hours (6 AM to 10 PM) are $\mathdollar 0.178$/kWh and off-peak hours are $\mathdollar 0.137$/kWh. Demand charge $K_{\mathrm{DC}}$ is $\mathdollar 11.67$/kW. The penalty for missing SoC is set at $K_{\mathrm{SOC}}=\mathdollar 0.5$/kWh, and battery degradation penalty is set to $K_{\mathrm{batt}}=\mathdollar 0.05$/kWh for discharging. The external charging cost is set to $\bar{E}=\mathdollar 0.30$/kWh, reflecting the average residential EV charging rate in the area, including amortized charger installation costs. More details in the Appendix F.
@@ 6. Experiments
Baseline Policies. We evaluate our framework against a series of progressively sophisticated baselines detailed in Table 1. The analysis begins with the Building-Only reference, which establishes the ideal operational cost without any EV load. We then model the common real-world scenario of Uncoordinated Charging (al2021slow), where EVs charge for free upon arrival. Building upon this, the Smart Charging + Free Cost policy (liu2025) isolates the benefit of centralized coordination by introducing an intelligent scheduling algorithm ($\pi^{\text{ch}}$) while keeping charging free for users. To introduce a simple economic model, we evaluate the Smart Charging + Fixed Cost policy, which sets the user price based on the local utility’s peak energy rate (svp). Advancing this, the Smart Charging + Utility-based Cost policy represents a purely operator-centric model that pairs the scheduling algorithm with our utility based pricing function (Eq. (3)) to maximize building savings. Finally, to benchmark against state-of-the-art interactive systems, we include the Menu-based Negotiation policy (bae2021inducing; ghosh2017menu), which introduces user choice and the ability to reject offers, serving as a direct precursor to our proposed framework.
7. Results

### v3

@@ Abstract
Keywords
1 Introduction
Figure 1: The CONSENT framework structures EV charging as a simple five-step interaction between the user and the building’s energy management system. Users arrive, submit their baseline request, and the system’s key contribution: Incentive Generation creates price–flexibility options based on forecasted load and energy prices (Panel 3). These options offer monetary discounts in exchange for modest flexibility in SoC or departure time. The user selects one option, after which the MPC engine executes the negotiated plan while ensuring the vehicle meets the agreed SoC and supporting building-level peak shaving.
@@ 1 Introduction
The rapid growth of electric vehicles (EVs), fueled by sustainability goals and advances in battery technology [2, 28], presents a significant challenge for energy management. When EV charging is uncoordinated, it can trigger sharp demand spikes, inflate energy costs, and threaten grid stability [14]. This has spurred research into coordinated charging, particularly in the vehicle-to-building (V2B) context where operators can leverage parked EVs for peak shaving [27]. However, creating an effective V2B system requires navigating a set of deeply interconnected challenges that span both technical control and human behavior.
@@
At the core of the control problem lies the interplay between uncertainty and long-term rewards [13]. A charging controller must operate with incomplete knowledge of future building loads and EV arrivals, yet its decisions are judged against long-term objectives like minimizing the monthly peak demand charge, a sparse and delayed reward. This makes myopic, reactive control insufficient. This technical challenge is further complicated by the need for incentive alignment [5]. The operator’s goal of deferring charging to minimize electricity costs, driven by both energy prices and monthly peak-power demand charges, directly conflicts with the user’s desire for convenience and a full battery [10]. A negotiation must therefore quantify the value of a user’s flexibility, but this value is itself a function of an uncertain future. Finally, any negotiation is vulnerable to strategic misreporting, where users may exaggerate their needs to extract larger incentives, undermining the system’s integrity. Addressing these issues requires more than just efficient scheduling; it demands a strategic bridge between the building’s stationary energy system and the dynamic mobility needs of its users. While prior work addresses these challenges in isolation, a key frontier remains: the co-optimization of the real-time charging control policy with a strategic, user-facing negotiation policy. This requires a framework that both makes optimal control decisions and generates fair, personalized incentives. While methods like Reinforcement Learning (RL) have shown promise for handling uncertainty, the co-optimization task particularly benefits from approaches that can explicitly handle hard constraints and adapt to new scenarios without extensive re-training. For this reason, we posit that the explicit constraint handling and computational efficiency of Model Predictive Control (MPC) provide a more direct and scalable solution [20, 18, 15]. An MPC-based approach offers robust uncertainty handling through sampling, which we show is comparable to an RL baseline, establishing MPC as the ideal engine for our framework.
@@
Leveraging this MPC foundation, we introduce our negotiation framework, CONSENT, which directly addresses the shortcomings of prior work [13, 9, 5]. To achieve incentive alignment, users are presented with tailored incentives for voluntary adjustments to their requested state-of-charge (SoC) or departure time, grounded in empirical survey data. To prevent strategic misreporting, our mechanism includes formal guarantees of strategy-proofness, budget feasibility, and voluntary participation. By integrating this mechanism and validating with real-world data, we demonstrate that operator savings can be aligned with high user participation and satisfaction. Our contributions include:
@@
2 Related Research
@@
Research in EV charging systems has largely advanced along two parallel streams. The first focuses on the charging control problem, where the goal is to optimize charging schedules to minimize operator costs like peak demand [10, 7]. A wide array of methods have been developed, including rule-based heuristics and mathematical programming [17, 23, 25]. More recent advances use Model Predictive Control (MPC) for real-time optimization, effectively adapting to dynamic loads, renewables, and prices [8, 29]. However, this entire stream of research, including stochastic MPC [20], typically solves the problem from the operator’s perspective by assuming full system control, thereby overlooking the critical challenge of securing voluntary user participation [26, 19].
@@
Recognizing that user compliance cannot be guaranteed, the second stream of research focuses on the socio-technical challenge of user negotiation and cooperation. This work has produced incentive-based frameworks like menu-based pricing [9] and iterative negotiation [27, 12] to provide users with choices. These are often informed by behavioral models that predict user responses to inconvenience and rewards [21], with a key finding being the importance of user heterogeneity [4, 31]. Despite being user-centric, these frameworks are often designed for large-scale aggregators and rarely integrated with the specific, real-time operational constraints of a single building.
@@
The limitations of these separate streams highlight a significant gap at their intersection: the co-optimization problem, which requires strategically coupling real-time control with empirically-grounded, user-centric negotiation. Tackling this frontier requires advanced methods capable of handling uncertainty and complex decisions. Reinforcement Learning (RL) has emerged as a powerful tool for learning control policies in such stochastic environments [13, 19]. However, for the specific co-optimization challenge, Model Predictive Control (MPC) presents a compelling alternative. While RL excels in exploration, the co-optimization problem benefits from MPC’s ability to explicitly handle hard constraints (e.g., user choices, charger limits) and efficiently evaluate hypothetical negotiation scenarios. By using Monte-Carlo sampling, MPC can robustly manage uncertainty, establishing it as a powerful engine for a framework that must bridge control and negotiation. Accordingly, our work addresses this co-optimization gap. We develop an MPC-based negotiation framework that operates directly on personalized user requests and is calibrated with empirical survey data, enabling voluntary and incentive-compatible cooperation at the building-user interface.
3 Problem Formulation
Figure 2: Negotiation workflow: Upon arrival, users receive forecast-based offers. Accepted offers are then optimally scheduled to meet the negotiated SoC.
@@ 3.1 Specifications
$\begin{split}&J^{\pi}(\bm{\theta}\mid\mathcal{W})=\sum_{t\in\mathcal{T}}g_{t}+K_{\mathbf{DC}}\cdot p^{\max}\\ &+K_{\mathbf{SOC}}\sum_{v\in\mathcal{V}}\left(e^{v}_{\mathrm{req}}-e^{v}_{\mathrm{dep}}\right)+K_{\mathbf{batt}}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q_{t}^{v}\end{split}$ (1)
@@
$\begin{split}U^{\bm{\theta}_{\setminus v},\pi^{*}}(\theta_{v})=\mathbb{E}_{\bm{\hat{\theta}}_{\setminus v}\sim\psi}\Big[J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}\mid\mathcal{W}^{\prime})\\ -J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v}\mid\mathcal{W}^{\prime})\Big]\end{split}$ (2)
@@ 4 Approach
4.1 EV-Charger Assignment & Charging Policy
@@
4.2 Formulating the User Negotiation Options
@@
As EVs depart on schedule, chargers free up, and the system state updates to reflect the connected EVs and their SoCs, charger availability, building load, and any new arrivals, enabling accurate tracking of charging and user decisions. The SMDP state transition kernel $\mathcal{P}(s_{k+1},\Delta t_{k}\mid s_{k},a_{k};\pi^{ch})$ combines: (i) the stochastic EV inter-arrival distribution; and (ii) the deterministic evolution under charging policy $\pi^{ch}$ during $\Delta t_{k}$. We do not explicitly define the state-action transition function, as our simulator leverages data-driven generative models trained on actual operational records [24].
@@ 4.3 Negotiation Policy
$\zeta^{v}=\sum_{t\in\mathcal{T}^{v}}g^{v}_{t}w_{t}-\alpha\cdot U({\theta}_{v})$ (3)
@@
Our negotiation engine selects incentive-backed charging offers for each arriving by integrating a Monte Carlo sampling based model predictive control (MC-MPC). Our approach belongs to the family of stochastic Model Predictive Control (SMPC) methods [16]. Algorithm 1 presents the user negotiation process.
@@
For each flexibility level $l$, the controller explores the corresponding continuous action subset $\mathcal{A}_{l}$ (from the SMDP) and solves a Mixed Integer Linear Program (MILP) implementing a Sample Average Approximation  [20]. The MILP minimizes the average total cost over the sampled scenarios $\mathcal{F}$, yielding an optimal candidate action $a_{l}^{*}$ for each $l$:
$a^{*}_{l}=\min_{\mathcal{A}_{l}}$ $\frac{1}{|\mathcal{F}|}\sum_{f\in\mathcal{F}}\Bigg[\sum_{i=0}^{H(t)}g_{i,f}+K_{\mathrm{DC}}\cdot p^{\max}_{f}$ (4)
@@ 20  Update EV $v_{k}$’s requirements based on the negotiated choice $\theta_{k}$;
5 Survey and User Modeling
@@
User acceptance during negotiation is modeled via a logit-based choice function [5]. The utility of acceptance for option $l$ by user type $n$ for charging cost $\zeta_{l}$ and given an external charging cost $\bar{E}$, is
$Y^{n}_{l}=\Big[\bar{E}-(\zeta_{l}+I^{n}_{l})\Big]$ (5)
@@
Policy Description Building Cost User Cost Monthly User Cost Reject (%)
(\$/month) (\$/kWh) (\$/month)
@@ 1 Building-Only Reference $8484\pm 2185$
2 Uncoordinated Charging [3] + Free Cost $9005\pm 2184$ $0.00$
3 Smart Charging + Free Cost [13] $8785\pm 2144$ $0.00$
4 Smart charging + Fixed User Cost [22] $8457\pm 2144$ $0.178$ $328\pm 61$
@@
6 Menu-based Negotiation [5, 9] $8479\pm 2153$ $0.157$ $309\pm 63$ 34.89
@@
Table 1: Comparative Evaluation of V2B Policies. The table progresses from a building-only reference to increasingly sophisticated baselines, culminating in our proposed negotiation framework (CONSENT). Lower values are better for all the metrics.
6 Experiments
We evaluate our negotiation framework using a modular simulator [24] of a building with multiple chargers, dynamic energy and demand charges, and realistic EV arrival and departure patterns. It captures stochastic user preferences, building load forecasts, and charger constraints, supporting online or ahead-of-time scheduling.
Hardware. Experiments ran on a 32-core, 6.2 GHz, 128 GB RAM machine using IBM CPLEX [11] for MILP.
Data Collection. We use data from a commercial site in Santa Clara, California (May 2023 – November 2024), including building power usage, charger activity, and vehicle telemetry. All EVs used support bidirectional charging, and charging is modeled with a three-segment piecewise linear SoC curve. We sampled $100$ such episodes, evenly split into $50$ for future sampling and $50$ for testing. Only weekdays are analyzed, as employee presence and demand charges are negligible on weekends. Figures 8 and 9 (Appendix F) show EV arrival/departure and SoC distributions. The site has controllable, 10 unidirectional (0–20 kW) and 5 bidirectional (±20 kW) chargers. Silicon Valley Power’s time-of-use and demand charge rates [22] define $\mathcal{W}$ where peak hours (6 AM to 10 PM) are $\mathdollar 0.178$/kWh and off-peak hours are $\mathdollar 0.137$/kWh. Demand charge $K_{\mathrm{DC}}$ is $\mathdollar 11.67$/kW. The penalty for missing SoC is set at $K_{\mathrm{SOC}}=\mathdollar 0.5$/kWh, and battery degradation penalty is set to $K_{\mathrm{batt}}=\mathdollar 0.05$/kWh for discharging. The external charging cost is set to $\bar{E}=\mathdollar 0.30$/kWh, reflecting the average residential EV charging rate in the area, including amortized charger installation costs. More details in the Appendix F.
@@ 6 Experiments
Baseline Policies. We evaluate our framework against a series of progressively sophisticated baselines detailed in Table 1. The analysis begins with the Building-Only reference, which establishes the ideal operational cost without any EV load. We then model the common real-world scenario of Uncoordinated Charging [3], where EVs charge for free upon arrival. Building upon this, the Smart Charging + Free Cost policy [13] isolates the benefit of centralized coordination by introducing an intelligent scheduling algorithm ($\pi^{\text{ch}}$) while keeping charging free for users. To introduce a simple economic model, we evaluate the Smart Charging + Fixed Cost policy, which sets the user price based on the local utility’s peak energy rate [22]. Advancing this, the Smart Charging + Utility-based Cost policy represents a purely operator-centric model that pairs the scheduling algorithm with our utility based pricing function (Eq. (3)) to maximize building savings. Finally, to benchmark against state-of-the-art interactive systems, we include the Menu-based Negotiation policy [5, 9], which introduces user choice and the ability to reject offers, serving as a direct precursor to our proposed framework.
7 Results